cos A+cos 2 A=1 , Then sin 2 A+Sin 4 A=?

cos A+cos2 A=1 , Then sin2A+Sin4A=?

Grade:

1 Answers

Y RAJYALAKSHMI
9 years ago
cosA + cos2A = 1 => cos2A = 1 – cosA
=> sin2A = 1 – cos2A = 1 – (1 – cosA) = cosA
Now,  sin2A + sin4A = sin2A (1 + sin2A) = cosA ( 1 + cosA) = cosA + cos2A = 1
 
Ans:  1